# Xing Huang a), Feng-Yu Wang a)

### v2

@@ 4. $(B_{3})$
(3.3) $\begin{split}I_{t}^{f}(\mu,\phi)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ \zeta_{s}&:=\sigma_{s}^{*}(\sigma_{s}\sigma_{s}^{*})^{-1},\ \ \ s\in[0,t],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.\end{split}$
By $(B_{1})$ and (2.14), we find a constant $c>0$ such that
(3.4) $|I_{t}^{f}(\mu,\phi)|\leq\frac{c}{\sqrt{t}}\big{(}P_{t}|f|^{k^{*}}(\mu)\big{)}^{\frac{1}{k^{*}}}\|\phi\|_{L^{k}(\mu)},\ \ \mu\in\mathscr{P}_{k},\phi\in T_{\mu,k}.$
@@ 1. $(1)$
  2. $(2)$
    The limit
    (3.5) $\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon},\ \ \ \ t\in[0,T]$
    exists in $L^{k}(\Omega\rightarrow C([0,T],\mathbb{R}^{d}),\mathbb{P})$, and there exists a constant $c>0$ such that
    $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\nabla_{\phi}X_{t}^{\mu}|^{k}\Big{]}\leq c\|\phi\|_{L^{k}(\mu)}^{k},\ \ \mu\in\mathscr{P}_{k}.$
  3. $(3)$
    For any $t\in(0,T]$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, $P_{t}f$ is intrinsically differentiable on $\mathscr{P}_{k}$. Moreover, for any $\mu\in\mathscr{P}_{k}$ and $\phi\in T_{\mu,k}$,
    (3.6) $\begin{split}&D_{\phi}^{I}P_{t}f(\mu)=I_{t}^{f}(\mu,\phi)+\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &N_{s}:=\big{\{}\nabla^{\mathbb{B}}_{D_{\phi}^{I}P_{s}V(\mu)}B_{s}(X_{s}^{\mu},P_{s}^{\ast}\mu,\cdot)\big{\}}(P_{s}V(\mu)),\\ &\tilde{N}_{s}:=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{\ast}\mu)(X_{s}^{\mu}),\nabla_{\phi}X_{s}^{\mu}\big{\rangle}\big{]}|_{y=X_{s}^{\mu}},\end{split}$
[…]
By (3.4)-(3.6) and ($B_{3}$), we find a constant $c>0$ such that
$\|D^{I}P_{t}f(\mu)\|_{L^{k^{*}}(\mu)}\leq c\frac{\{P_{t}|f|^{k^{*}}(\mu)\}^{\frac{1}{k^{*}}}}{\sqrt{t}},\ \ t\in(0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k}.$
Remark 3.3.
Note that if $\nabla V$ is continuous and bounded, then the chain rule (see for instance [19, Proposition 3.1]) implies
(3.7) $D^{I}_{\phi}P_{t}V(\mu)=\mathbb{E}\nabla_{\nabla_{\phi}X_{t}^{\mu}}V(X_{t}^{\mu}).$
However, in the present case, $V$ only satisfies $(B_{2})$ so that (3.7) does not hold and there is essential difficulty in the study of $D^{I}_{\phi}P_{t}V$.
To explain the main idea in the proof of Theorem 3.2, we first figure out a sketch. By the definition of the intrinsic derivative, we intend to calculate
@@ 3. $(3)$
To this end, for any $\gamma\in\mathscr{P}_{k},$ consider the decoupled SDE (3.2) for $\gamma$ replacing $\mu$, i.e.
(3.9) $\begin{split}\text{\rm{d}}X_{t}^{x,\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{x,\gamma})+&B_{t}\big{(}X_{t}^{x,\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{x,\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],X_{0}^{x,\gamma}=x\in\mathbb{R}^{d}.\end{split}$
Let
$P_{t}^{\gamma}f(x):=\mathbb{E}[f(X_{t}^{x,\gamma})],\ \ \ x\in\mathbb{R}^{d},$
$P_{t}^{\gamma}f(\tilde{\gamma}):=\int_{\mathbb{R}^{d}}P_{t}^{\gamma}f\text{\rm{d}}\tilde{\gamma},\ \ t\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}),\gamma,\tilde{\gamma}\in\mathscr{P}_{k}.$
For $\varepsilon\geq 0$, let $X_{t}^{\mu_{\varepsilon},\gamma}$ be the solution of (3.9) with initial value $X_{0}^{\varepsilon}$, i.e,
$\begin{split}\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\gamma}=\big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\gamma})+&B_{t}\big{(}X_{t}^{\mu_{\varepsilon},\gamma},P_{t}^{\ast}\gamma,P_{t}V(\gamma)\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\gamma})\text{\rm{d}}W_{t},\\ &t\in[0,T],\ X_{0}^{\mu_{\varepsilon},\gamma}=X_{0}^{\varepsilon}.\end{split}$
@@
$P_{t}f(\mu_{\varepsilon})=P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})=\mathbb{E}[f(X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}})],\ \ \varepsilon\geq 0,t\in[0,T],f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
@@
By Lemma 2.3(1), for any $t\in(0,T],\varepsilon\geq 0$ and $f\in\mathscr{B}_{b}(\mathbb{R}^{d})$, we have
(3.11) $\begin{split}&\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}P_{t}^{\mu}f(\mu_{\varepsilon}):=\lim_{r\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon+r})-P_{t}^{\mu}f(\mu_{\varepsilon})}{r}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})\frac{1}{t}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)\\ &=:I_{t}^{f}(\varepsilon,\mu,\phi),\ \ t\in(0,T].\end{split}$
@@
$\lim_{\varepsilon\rightarrow 0}J_{2}f(t,\varepsilon)=\mathbb{E}\bigg{[}f(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\big{\{}N_{s}+\tilde{N}_{s}\big{\}},\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}$
@@
for $N_{s},\tilde{N}_{s}$ defined in (3.6), which involves in $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$ defined in (3.5). Therefore, we will first study $D^{I}P_{t}V(\mu)$ and $\nabla_{\phi}X_{t}^{\mu}$.
To this end, let $u_{t}=u_{t}^{\lambda_{0},\mu},\Theta_{t}=\Theta_{t}^{\lambda_{0},\mu}$ be defined in (2.7) and (2.10) for large enough $\lambda_{0}>0$ such that (2.9) holds for $\lambda=\lambda_{0}$, and let
$\mathscr{C}_{k}:=\left\{\tilde{v}\in L^{k}\big{(}\Omega\rightarrow C([0,T];\mathbb{R}^{d}),\mathbb{P}\big{)}\ \text{is\ }\mathscr{F}_{t}\text{-adapted,\ }\tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu})\right\},$
which is complete under the metric
$\rho_{\delta}(\tilde{v}^{1},\tilde{v}^{2}):=\left(\mathbb{E}\sup_{t\in[0,T]}\text{\rm{e}}^{-\delta kt}|\tilde{v}_{t}^{1}-\tilde{v}^{2}_{t}|^{k}\right)^{\frac{1}{k}},\ \ \tilde{v}^{1},\tilde{v}^{2}\in\mathscr{C}_{k},\delta>0.$
For any $s\in[0,T]$, $\eta\in L^{k}(\Omega\rightarrow\mathbb{R}^{d},\mathbb{P})$, define
$\psi_{s}(v):=\mathbb{E}\big{[}\big{\langle}\{D^{L}B_{s}(y,\cdot,P_{s}V(\mu))\}(P_{s}^{*}\mu)(X_{s}^{\mu}),(\nabla\Theta_{s}(X_{s}^{\mu}))^{-1}\eta\big{\rangle}\big{]}|_{y=X_{s}^{\mu}}.$
@@
Recall that $\tilde{\alpha}$ is defined in (2.35). For any $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$, we denote
(3.13) $\begin{split}&N_{s}(v):=\nabla_{v_{s}}^{\mathbb{B}}B_{s}(X_{s}^{\mu},P_{s}^{*}\mu,\cdot)(P_{s}V(\mu)),\ \ \tilde{N}_{s}(\tilde{v}):=\psi_{s}(\tilde{v}_{s}),\end{split}$
and write
@@
$\begin{split}&\tilde{I}_{t}^{V}(\mu,v,\tilde{v})=\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v})+\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}),\\ &\tilde{I}_{t}^{V,1}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}\Big{\langle}\zeta_{s}(X_{s}^{\mu})N_{s}(v),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\\ &\tilde{I}_{t}^{V,2}(\mu,v,\tilde{v}):=\mathbb{E}\bigg{[}V(X_{t}^{\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{\mu})\tilde{N}_{s}(\tilde{v}),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]},\ \ t\in[0,T].\end{split}$
Consider the following equation for $(v,\tilde{v})\in C([0,T];\mathbb{B})\times\mathscr{C}_{k}$:
@@
(3.14) $\begin{split}&\text{\rm{d}}\tilde{v}_{t}=\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\big{\}}(X_{t}^{\mu})\Big{]}\text{\rm{d}}t\\ &+\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t+{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t\\ &+\Big{[}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\Big{]}\,\text{\rm{d}}W_{t},\ \ \tilde{v}_{0}=\nabla_{\phi(X_{0}^{\mu})}\Theta_{0}(X_{0}^{\mu}),\\ &v_{t}={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}},\ \ t\in[0,T].\end{split}$
In the following two subsections, we prove the well-posedness of (3.14) and the formulas
(3.15) $\begin{split}&D_{\phi}^{I}P_{t}V(\mu)={\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\,v_{t},\ \ t\in(0,T],\\ &\nabla_{\phi}X_{t}^{\mu}:=\lim_{\varepsilon\downarrow 0}\frac{X_{t}^{\mu_{\varepsilon}}-X_{t}^{\mu}}{\varepsilon}=(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t},\ \ \ t\in[0,T].\end{split}$
@@
(3.16) $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\|v_{t}(\mu,\phi)\|_{\mathbb{B}}\leq{c\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T],$
and for any $j\geq 1$, there exists a constant $\tilde{c}>0$ such that
(3.17) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}|\tilde{v}_{t}(\mu,\phi)|^{j}|\mathscr{F}_{0}\Big{]}\leq\tilde{c}|\phi(X_{0}^{\mu})|^{j}+{\tilde{c}\|\phi\|_{L^{k}(\mu)}^{j}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T].$
@@
By Jensen’s inequality, for estimate (3.17) we only need to consider $j\geq k.$
As in the proof of [29, Lemma 4.1], there exists a constant $\theta>1$ and $0<f\in\mathscr{B}([0,T]\times\mathbb{R}^{d})$ such that
(3.18) [$\big{\|}\nabla\big{[}\{\nabla\Theta\}\sigma\big{]}\big{\|}\leq f,\ \ \ \mathbb{E}\int_{0}^{T}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s<\infty.$
For any $n\geq 0$, define
$\tau_{n}:=T\wedge\inf\left\{t\geq 0:\int_{0}^{t}f_{s}(X_{s}^{\mu})^{2\theta}\text{\rm{d}}s\geq n\right\},$
$\mathscr{C}_{k,n}:=\big{\{}(z_{\cdot\wedge\tau_{n}}),\ \ z\in\mathscr{C}_{k}\big{\}}.$
We have $\lim_{n\rightarrow\infty}\tau_{n}=T$. Let
$\mathscr{V}_{0}:=\big{\{}v\in C([0,T];\mathbb{B}):\ v_{0}=0\big{\}}.$
Recall that $\tilde{\alpha}$ is defined in (2.35). By (3.18), $\|\nabla B\|_{\infty}<\infty$ due to (B), and
(3.19) $\int_{0}^{t}\frac{\alpha(rs^{\frac{1}{2}})^{2}}{s\tilde{\alpha}(rs^{\frac{1}{2}})}\text{\rm{d}}s=2\int_{0}^{rt^{\frac{1}{2}}}\frac{\alpha(s)^{2}}{s\tilde{\alpha}(s)}\text{\rm{d}}s=4\int_{0}^{rt^{\frac{1}{2}}}\tilde{\alpha}^{\prime}(s)\text{\rm{d}}s=4\tilde{\alpha}(rt^{\frac{1}{2}}),\ \ r\geq 0,$
we see that for any $(v,\tilde{v})\in\mathscr{V}_{0}\times\mathscr{C}_{k},$
$H_{s}(v,\tilde{v})$ $:=\tilde{v}_{0}+\int_{0}^{s}\nabla_{\psi_{t}(\tilde{v}_{t})}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}\Big{\{}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\{B_{t}(\cdot,P_{t}^{*}\mu,P_{t}V(\mu))+\lambda_{0}u_{t}\}\Big{\}}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}{\frac{\alpha(t^{\frac{1}{2}})}{\{\tilde{\alpha}(t^{\frac{1}{2}})t\}^{\frac{1}{2}}}}\nabla_{N_{t}(v)}\Theta_{t}(X_{t}^{\mu})\text{\rm{d}}t$
$+\int_{0}^{s}\nabla_{(\nabla\Theta_{t}(X_{t}^{\mu}))^{-1}\tilde{v}_{t}}\big{\{}(\nabla\Theta_{t})\sigma_{t}\big{\}}(X^{\mu}_{t})\,\text{\rm{d}}W_{t},\ \ s\in[0,T]$
is a continuous adapted process on $\mathbb{R}^{d}$ such that
$H_{\cdot\land\tau_{n}}:\mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{C}_{k,n},\ \ n\geq 1.$
On the other hand, let
(3.20) $\begin{split}&\bar{H}_{t}(v,\tilde{v}):={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{V}(\mu,\phi)+\tilde{I}_{t}^{V}(\mu,v,\tilde{v})\big{\}}.\end{split}$
We claim that it suffices to prove
@@
    for each $n\geq 1$, the map
    $\Phi^{(n)}:=(\bar{H},H_{\cdot\land\tau_{n}}):\ \mathscr{V}_{0}\times\mathscr{C}_{k,n}\rightarrow\mathscr{V}_{0}\times\mathscr{C}_{k,n}$
    is well-defined and has a unique fixed point $(v^{(n)},{\tilde{v}}^{(n)})$.

### v4

@@ 4. $(B_{3})$
(3.8) $\begin{split}I_{t}^{f}(\mu,\phi)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ &s\in[0,t],\mu\in\mathscr{P}_{k},\phi\in T_{\mu,k},f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).\end{split}$
By (B) and (3.3), we find a constant $c>0$ such that
(3.9) $|I_{t}^{f}(\mu,\phi)|\leq\frac{c}{\sqrt{t}}\big{(}P_{t}|f|^{k^{*}}(\mu)\big{)}^{\frac{1}{k^{*}}}\|\phi\|_{L^{k}(\mu)},\ \ \mu\in\mathscr{P}_{k},\phi\in T_{\mu,k},f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).$
@@ 1. $(1)$
    For any $t\in(0,T]$, $y\in\mathbb{R}^{d}$, $\nu\in\mathscr{P}_{k}$, $P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)$ is intrinsically differentiable on $\mu\in\mathscr{P}_{k}$, and there exists a constant $c>0$ such that
    $\sup_{(y,\nu)\in\mathbb{R}^{d}\times\mathscr{P}_{k}}\|D^{I}P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)](\mu)\|_{L^{\frac{k}{k-1}}(\mu)}\leq\frac{c\,\alpha(t^{\frac{1}{2}})}{\sqrt{t}},\ \ t\in(0,T],\mu\in\mathscr{P}_{k}.$
  2. $(2)$
    For any $t\in(0,T]$ and $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$, $P_{t}f$ is intrinsically differentiable on $\mathscr{P}_{k}$. Moreover, for any $\mu\in\mathscr{P}_{k}$ and $\phi\in T_{\mu,k}$,
    (3.10) $\begin{split}D_{\phi}^{I}P_{t}f(\mu)&=I_{t}^{f}(\mu,\phi)\\ &+\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x),\\ &N_{s}(\mu,\phi):=\big{\{}D_{\phi}^{I}P_{s}[D^{E}b_{s}^{(1)}(y,\nu)(\cdot)](\mu)\big{\}}|_{y=X_{s}^{x,\mu},\nu=P_{s}^{\ast}\mu},\end{split}$
@@ 2. $(2)$
By (3.9)-(3.10), we find a constant $c>0$ such that
$\|D^{I}P_{t}f(\mu)\|_{L^{k^{*}}(\mu)}\leq c\frac{\{P_{t}|f|^{k^{*}}(\mu)\}^{\frac{1}{k^{*}}}}{\sqrt{t}},\ \ t\in(0,T],f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}),\mu\in\mathscr{P}_{k}.$
To explain the main idea in the proof of Theorem 3.3, we first figure out a sketch. By the definition of the intrinsic derivative, we intend to calculate for any $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$,
@@
To this end, for any $\mu\in\mathscr{P}_{k},x\in\mathbb{R}^{d}$, recall that $X_{t}^{x,\mu}$ solves the decoupled SDE (3.2), and
$P_{t}^{\mu}f(x)=\mathbb{E}[f(X_{t}^{x,\mu})],\ \ \ x\in\mathbb{R}^{d}.$
Define
$P_{t}^{\mu}f(\tilde{\mu}):=\int_{\mathbb{R}^{d}}P_{t}^{\mu}f\text{\rm{d}}\tilde{\mu},\ \ t\geq 0,f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}),\mu,\tilde{\mu}\in\mathscr{P}_{k}.$
For $\varepsilon\geq 0$, let $X_{t}^{\mu_{\varepsilon},\mu}$ be the solution of (3.2) with initial value $X_{0}^{\varepsilon}$, i.e,
$\begin{split}\text{\rm{d}}X_{t}^{\mu_{\varepsilon},\mu}=\big{\{}b_{t}^{(0)}(X_{t}^{\mu_{\varepsilon},\mu})+&b_{t}^{(1)}\big{(}X_{t}^{\mu_{\varepsilon},\mu},P_{t}^{\ast}\mu\big{)}\big{\}}\text{\rm{d}}t+\sigma_{t}(X_{t}^{\mu_{\varepsilon},\mu})\text{\rm{d}}W_{t},\\ &t\in[0,T],\ X_{0}^{\mu_{\varepsilon},\mu}=X_{0}^{\varepsilon}.\end{split}$
@@
$P_{t}f(\mu_{\varepsilon})=P_{t}^{\mu_{\varepsilon}}f(\mu_{\varepsilon})=\mathbb{E}[f(X_{t}^{\mu_{\varepsilon},\mu_{\varepsilon}})],\ \ \varepsilon\geq 0,t\in[0,T],f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d}).$
@@
By Lemma 3.2, for any $t\in(0,T],\varepsilon\geq 0$ and $f\in\mathscr{B}_{k-1,b}(\mathbb{R}^{d})$, we have
$\begin{split}&\frac{\text{\rm{d}}}{\text{\rm{d}}\varepsilon}P_{t}^{\mu}f(\mu_{\varepsilon}):=\lim_{r\downarrow 0}\frac{P_{t}^{\mu}f(\mu_{\varepsilon+r})-P_{t}^{\mu}f(\mu_{\varepsilon})}{r}\\ &=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x+\varepsilon\phi(x),\mu})\frac{1}{t}\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x+\varepsilon\phi(x),\mu})\nabla_{\phi(x)}X_{s}^{x+\varepsilon\phi(x),\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
@@
$\lim_{\varepsilon\rightarrow 0}J_{2}f(t,\varepsilon)=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}f(X_{t}^{x,\mu})\int_{0}^{t}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})N_{s}(\mu,\phi),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x)$
@@
for $N_{s}(\mu,\phi)$ defined in (3.10), which involves in $D_{\phi}^{I}\{P_{s}[D^{E}b_{s}^{(1)}(y,\nu)(\cdot)](\mu)\}$. Therefore, we will first study $D_{\phi}^{I}\{P_{s}[D^{E}b_{s}^{(1)}(y,\nu)(\cdot)](\mu)\}$.
@@
Recall that $\tilde{\alpha}$ is defined in (2.40). For any $V\in\mathscr{B}_{b}([0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k};\mathbb{R}^{d})$, the set of bounded and measurable $\mathbb{R}^{d}$-valued functions on $[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}$ and $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, we write
@@
(3.14) $\begin{split}\tilde{I}_{t}^{V}(y,\nu)&:=\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}D^{E}b_{t}^{(1)}(y,\nu)(X_{t}^{x,\mu})\\ &\qquad\qquad\quad\times\int_{0}^{t}\frac{\alpha(s^{\frac{1}{2}})}{\{\tilde{\alpha}(s^{\frac{1}{2}})s\}^{\frac{1}{2}}}\Big{\langle}\zeta_{s}(X_{s}^{x,\mu})V_{s}(X_{s}^{x,\mu},P_{s}^{\ast}\mu),\ \text{\rm{d}}W_{s}\Big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
By $(B_{2})$, for any $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, we have
$|D^{E}b_{t}^{(1)}(y,\nu)(\cdot)|$ $\leq\alpha(|\cdot|)+|D^{E}b_{t}^{(1)}(y,\nu)(0)|$
$\leq c_{0}(1+|\cdot|^{k-1})+|D^{E}b_{t}^{(1)}(y,\nu)(0)|.$
So, $I_{t}^{f}(\mu,\phi)$ for $f=D^{E}b_{t}^{(1)}(y,\nu)(\cdot)$ in (3.8) is well-defined, and we denote
(3.15) $\begin{split}I_{t}^{\mu,\phi}(y,\nu)&:=\frac{1}{t}\int_{\mathbb{R}^{d}}\mathbb{E}\bigg{[}D^{E}b_{t}^{(1)}(y,\nu)(X_{t}^{x,\mu})\\ &\qquad\quad\times\int_{0}^{t}\big{\langle}\zeta_{s}(X_{s}^{x,\mu})\nabla_{\phi(x)}X_{s}^{x,\mu},\text{\rm{d}}W_{s}\big{\rangle}\bigg{]}\mu(\text{\rm{d}}x).\end{split}$
Consider the following equation for $V\in\mathscr{B}_{b}([0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k};\mathbb{R}^{d})$:
(3.16) $\begin{split}&V_{t}(y,\nu)={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{\mu,\phi}(y,\nu)+\tilde{I}_{t}^{V}(y,\nu)\big{\}},\ \ t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
If this equation has a unique solution, we denote it by $V_{t}(y,\nu)=v_{t}^{\mu,\phi}(y,\nu)$ for $(t,y,\nu)\in[0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k}$, to emphasize the dependence on $\mu$ and $\phi$.
In the following two subsections, we prove the well-posedness of (3.16) and establish the formula
@@
(3.17) $\begin{split}&D_{\phi}^{I}\{P_{t}[D^{E}b_{t}^{(1)}(y,\nu)(\cdot)]\}(\mu)={\frac{\alpha(t^{\frac{1}{2}})}{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}}\,v_{t}^{\mu,\phi}(y,\nu),\ \ t\in(0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}.\end{split}$
@@
(3.18) $\sup_{\|\phi\|_{L^{k}(\mu)}\leq 1}\sup_{y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}}|v_{t}^{\mu,\phi}(y,\nu)|\leq{c\,\sqrt{\tilde{\alpha}(t^{\frac{1}{2}})}},\ \ \mu\in\mathscr{P}_{k},t\in[0,T].$
@@
Let
$\mathscr{V}_{0}:=\big{\{}V\in\mathscr{B}_{b}([0,T]\times\mathbb{R}^{d}\times\mathscr{P}_{k};\mathbb{R}^{d}):\ V_{0}=0\big{\}},$
which is a Banach space under the uniform norm. For $V\in\mathscr{V}_{0},$ let
$\|V_{t}\|_{\infty}=\sup_{y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}}|V_{t}(y,\nu)|,\ \ t\in[0,T]$
and for any $t\in[0,T],y\in\mathbb{R}^{d},\nu\in\mathscr{P}_{k}$, let
(3.19) $\begin{split}&\{H(V)\}_{t}(y,\nu):={\frac{\{t\tilde{\alpha}(t^{\frac{1}{2}})\}^{\frac{1}{2}}}{\alpha(t^{\frac{1}{2}})}}\big{\{}I_{t}^{\mu,\phi}(y,\nu)+\tilde{I}_{t}^{V}(y,\nu)\big{\}}.\end{split}$
Then it suffices to prove
@@
[…]